# Equilibria in non-Euclidean geometries

### v1

@@ 1. Introduction
Centroids and static equilibrium points of convex bodies have been in the focus of research since the work of Archimedes, whose results have been used in naval design even in the 18th century [23]. These concepts appear not only in engineering see e.g. [21, 38], but also in geology [13, 27], biology [14, 8], medicine [1] and in many other disciplines. Among the classical problems in pure mathematics related to these concepts we can mention problems related to floating bodies ( see e.g. [32, 35, 26]), Grünbaum’s inequality [22] and its variants, and the Busemann-Petty centroid inequality [33] and its variants. There are also many mathematical results related to the number of equilibrium points of convex bodies. As examples, we may mention the construction of a double-tipping tetrahedron by Heppes [24], or the monostable polyhedron of Conway and Guy in [20]. Recent solutions of two problems on monostable polyhedra [29] by the first-named author, proposed by Conway and Guy [20] in 1969 are also worth mentioning here. As for general convex bodies, a result of Domokos, Papadopoulos and Ruina [12] states that a nondegenerate convex body in the Euclidean plane has at least four equilibria, whereas the construction of a mono-monostatic convex body in [39] by Várkonyi and Domokos, called Gömböc, shows that the same statement does not hold for $3$-dimensional convex bodies. We note that the latter body is well known outside the mathematical community and has found applications in many other disciplines [^1] For example, it was chosen to symbolize monetary stability by the National Bank of Hungary, and also as the insignia for the Steven Smale Prize by the Society for the Foundations of Computational Mathematics. For more information on the mathematical aspects of centroids and equilibria, the interested reader is referred to the recent survey [30].
@@
In the past century, numerous attempts have been made to generalize the notions of centroids to non-Euclidean geometries, mostly involving spherical and hyperbolic spaces. In particular, centroids of spherical triangles have been investigated by Fog [18] and Fabricius-Bjerre [16] in the 1940s. Their definition uses the model of spherical space as a hypersurface embedded in a Euclidean space (for such an approach, see also [7]). In 1993, Gal’perin [19] gave a thorough investigation of possible definitions of centroids of systems of points in spaces of constant curvature, most importantly in spherical and hyperbolic space. His results were recently generalized for $k$-dimensional manifolds in [9]. A recent paper of Besau et al. [5] independently defines the centroid of a spherical set in the same way.
@@
The main goal of this paper is to extend the above investigation to equilibrium properties of convex bodies in non-Euclidean spaces, more specifically in spherical, hyperbolic and normed spaces. In particular, we investigate how the results in [12] and [39] can be generalized for convex bodies in these spaces. During this investigation, we always imagine a $d$-dimensional normed space to be equipped with an underlying Euclidean space, and, for $d\geq 3$, call it rotationally symmetric if there is a (d2)(d-2)-dimensional linear subspace $L$ in it such that the norm is invariant under any rotation around $L$.
@@
$d_{H}(A,B)=\inf\{d\geq 0:\forall a\in A\exists b\in B\hbox{ with }d(a,b)\leq d,\\ \hbox{ and }\forall b\in B\exists a\in A\hbox{ with }d(a,b)\leq d\}.$
@@ 2.1. Centroids in non-Euclidean spaces
where $d(H,x)$ denotes the signed Euclidean distance of $H$ and $x$; the sign is determined by the orientation of $H$. This formula is a consequence of the Euclidean rule of the lever for convex bodies, and the integral in (2.2), denoted by $M_{H}(K)$, is called the first moment of $K$ with respect to $H$.
@@
where $d(H,x)$ denotes the signed spherical distance of $H$ and $x$; the sign is determined by the orientation of $H$. Here the integral in (2.4), denoted by $M_{H}(K)$, is called the first moment of $K$ with respect to $H$.
@@
where $d_{H}^{e}(H,x)$ denotes the signed hyperbolic distance of $H$ and $x$; the sign is determined by the orientation of $H$. Here the integral in (2.6), denoted by $M_{H}(K)$, is called the first moment of $K$ with respect to $H$.
@@
    By a classical theorem of Haar, every meaningful definition of volume in $\mathbb{M}^{d}$ is a scalar multiple of $d$-dimensional Lebesgue measure (see e.g. [2, 4, 28]).
@@ 2.2. Equilibria in non-Euclidean spaces
Consider a convex body $K$ in the Euclidean space $\mathbb{R}^{d}$. If the point $q\in\operatorname{bd}(K)$ has the property that the hyperplane through $q$ and orthogonal to qc(K)q-c(K) supports $K$, then we say that $q$ is an equilibrium point of $K$. It is well known (see e.g. [10, 11, 29]) that if $K$ is smooth, then the equilibrium points of $K$ coincide with the critical points of the Euclidean distance function $x\mapsto d(x,c(K))$, measured from $c(K)$ and restricted to $\operatorname{bd}(K)$. We remark that a convex body $K\subset\mathbb{R}^{3}$ is called smooth if for any boundary point $x$ of $K$ there is a unique supporting hyperplane of $K$ at $q$; this property coincides with the property that $\operatorname{bd}(K)$ is a $C^{1}$-class submanifold of $\mathbb{R}^{d}$ (cf. [36]).
@@
Let $K$ be a nondegenerate convex body $K$ in $\mathbb{R}^{d}$ with $d=2$ or $d=3$.
@@
By Corollary 2.8, to determine the numbers of the different types of equilibrium points of a nondegenerate convex body in $\mathbb{R}^{2}$, it is sufficient to know the number $S$ of its stable points, and in $\mathbb{R}^{3}$ the numbers of its stable and unstable points. Thus, every nondegenerate convex body in $\mathbb{R}^{2}$ belongs to a class $\{S\}_{E}$ consisting of the bodies having $S$ stable and $S$ unstable points, and in $\mathbb{R}^{3}$ to a class $\{S,U\}_{E}$ consisting of the bodies having $S$ stable, $U$ unstable and S+U2S+U-2 saddle-type points. We remark that the global minimizers and maximizers of the Euclidean distance function are stable and unstable points of the body $K$, and thus, the quantities $S,U$ above are positive integers.
@@
For a convex body $K$ in a normed space we follow the same approach. More specifically, if $K$ is a convexin the normed space $\mathbb{M}^{d}$ and $q\in\operatorname{bd}(K)$, we say that $q$ is an equilibrium point of $K$ if there is a hyperplane $H$ supporting $K$ at $q$ that is Birkhoff orthogonal to the segment $[c(K),q]$. We recall that this is equivalent to the property that a homothetic copy of the unit ball $M$ of $\mathbb{M}^{d}$, centered at $c(K)$, touches $K$. Thus, if both $K$ and $M$ are smooth, then the equilibrium points of $K$ coincide with the critical points of the normed distance function $x\mapsto d_{M}(c(K),x)$ restricted to $\operatorname{bd}(K)$. In the paper we discuss the equilibrium points of convex bodies in planes with strictly convex, smooth norms: in this case for any hyperplane $H$ and point $x\notin H$ there is a unique segment $[x,q]$ with $q\in H$ to which it is Birkoff orthogonal, and for every line $[x,q]$ there is a unique hyperplane through $q$ that is Birkhoff orthogonal to it.
@@
Assume that $x$ is a local minimum point of $\rho_{K}$. Then, within a neighborhood $V$ of the corresponding point $q_{x}$ of $\operatorname{bd}(K)$, $K$ contains the set $V\cap D$ centered at $c(K)$ and radius $\rho_{K}(x)$. Since $D$ is smooth and $K$ is convex, the unique supporting line of $D$ at $q_{x}$ supports $K$ as well. Thus, $q_{x}$ is an equilibrium point of $K$. It follows similarly that for every local maximum point $x$ of $\rho_{K}$, the corresponding point $q_{x}$ of $\operatorname{bd}(K)$ is an equilibrium point. Hence, it follows that $\rho_{K}$ is not constant on any nondegenerate arc of $S$, as otherwise $K$ has infinitely many equilibrium points.
@@ 4.1. The case $\mathbb{V}^{3}=\mathbb{S}^{3}$
Consider the unit sphere $\mathbb{S}^{3}$ of $\mathbb{R}^{4}$ centered at the origin $o$. Set the point $v_{N}=(0,0,0,1)$; we call this point the North Pole of $\mathbb{S}^{3}$, and the set of points of $\mathbb{S}^{3}$ with positive $x_{4}$-coordinates the Northern Hemisphere $N^{3}$ of $\mathbb{S}^{3}$. Let $H_{N}$ be the hyperplane of $\mathbb{R}^{4}$ with equation $\{x_{4}=1\}$. The gnomonic projection of $\mathbb{S}^{3}$ onto $H_{N}$ is defined as the map $g_{N}:N^{3}\to H_{N}$, where the image $g_{N}(x)$ of $x\in N^{3}$ is the intersection of $H_{N}$ and the halfline $\{\lambda x:\lambda>0\}$. Note that $g_{N}$ is a bijection between $N^{3}$ and $H_{N}$ with the property that $X\subset N^{3}$ is a greatcircle arc in $N^{3}$ iff its image $g_{N}(X)$ is a Euclidean segment in $H_{N}$. Thus, for any spherically convex body $K\subset N^{3}$, $g_{N}(K)$ is a convex body in $H_{N}$, regarded as a $3$-dimensional Euclidean space, and vice versa: if $L$ is a (Euclidean) convex body in $H_{N}$, then $g^{-1}(L)$ is a (spherical) convex body in $N^{3}$.
@@
$g_{N}^{-1}(x)=\left(\frac{x_{1}}{\sqrt{1+|x|_{N}^{2}}},\frac{x_{2}}{\sqrt{1+|x|_{N}^{2}}},\frac{x_{3}}{\sqrt{1+|x|_{N}^{2}}},\frac{1}{\sqrt{1+|x|_{N}^{2}}}\right)$
@@
For any $0<c\leq 1$, we have $f_{c}\left(\frac{\pi}{2}\right)=g_{c}\left(\frac{\pi}{2}\right)=\pi/2$, $f_{c}\left(-\frac{\pi}{2}\right)=g_{c}\left(-\frac{\pi}{2}\right)=-\pi/2$.
@@
(4.4) $a_{c}(\theta,\varphi)=\frac{\cos^{2}\varphi\cos^{2}f_{c}(\theta)}{\cos^{2}\frac{3\varphi}{2}\cos^{2}f_{c}(\theta)+\sin^{2}\varphi\cos^{2}g_{c}(\theta)},$
as illustrated by Fig. 2. We observe that if $\cos\frac{m\varphi}{2}\neq 0$ (i.e. if $\varphi$ is not an integer multiple of $\frac{2\pi}{m}$), then $a_{c}(\theta,\varphi)=\frac{1}{1+\tan^{2}\varphi\frac{\cos^{2}g_{c}(\theta)}{\cos^{2}f_{c}(\theta)}}$.
Figure 2. Diagram of $a_{c}(\theta,\varphi)$ versus $\varphi$ for two different values of $c$. If $c=1$ then $a_{c}=\cos(2\varphi)$ for all values of $\theta$ (left panel), whereas for other values of $c$ and $\theta\neq 0$, it becomes somewhat distorted. The right panel shows the values of $a_{c}(\theta,\varphi)$ for $c=0.1$, with $\theta=-0.45\pi$ (continuous line), $\theta=-0.225\theta$ (long dashed line), $\theta=0$ (dashed line), $\theta=0.225\pi$ (dashed-dotted line), and $\theta=0.45\pi$ (dotted line).
@@
and remark that by definition, for any value of the parameters we have $0\leq a_{c}(\theta,\varphi)\leq 1$. Thus, $\rho_{c}(u)$ is a convex combination of $f_{c}(\theta)$ and $g_{c}(\theta)$. This, by Lemma 4.3, implies that $1\leq\rho_{c}(u)\leq 1$ and $R(1-d)\leq R_{c,d}(u)\leq R(1+d)$. Thus, $K(c,d)$ exists for all $0\leq d<1$.
@@
Next, we prove that the properties described in (A)-(B) are satisfied for $K(c,d)$ for all $0<d<1$ and $0<c\leq 1$. To show them, we recall [11, Lemmas 3, 4]. We note that we state them using the notation introduced in this subsection.
@@
Now we prove that suitable elements of $\mathcal{F}_{m}$ satisfy the properties (C)-(E). For this purpose, we set some arbitrary small positive constant $\varepsilon>0$, and observe that the definition of $\rho_{c}$ implies that $\max\{\rho_{c}(u):u\in S\}=1$, and hence, if $0<d<\varepsilon$, then $d_{H}(K_{n}(c,d),RB)<R\varepsilon$. Thus, the assertion of Theorem 1.2 for $\mathbb{S}^{3}$ readily follows from Lemmas 4.7 and 4.8.
@@
There exist constants $0<c_{1}<c_{2}<1$, $0<d_{0}<1$ and a function $F:(0,d_{0})\to[c_{1},c_{2}]$ such that for any $d\in(0,d_{0})$ the spherical centroid of $K(F(d),d)$ is $v_{N}$.
@@
We note that the integral in (4.9) is also computed in [11], where the authors incorrectly stated that its value is $2.3168\ldots$; nevertheless, this typo in [11] does not influence the correctness of their proof. Next, we deal with Property (C). To do it, following [36] we denote by $\mathcal{C}^{2}_{+}$ the family of convex bodies whose boundary is $C^{2}$-class differentiable at every point, and has strictly positive Gaussian curvature everywhere.

### v2

@@ 1. Introduction
Centroids and static equilibrium points of convex bodies have been in the focus of research since the work of Archimedes, whose results were used in naval design even in the 18th century [23]. These concepts appear not only in physics, but also in engineering [21, 39], geology [13, 28], biology [14, 8], medicine [1] and in many other disciplines. Among the classical problems in pure mathematics related to these concepts we can mention problems related to floating bodies ( see e.g. [33, 36, 26]), Grünbaum’s inequality [22] and its variants, and the Busemann-Petty centroid inequality [34] and its variants. There are also many mathematical results related to the number of equilibrium points of convex bodies. As examples, we may mention the construction of a double-tipping tetrahedron by Heppes [24], or the monostable polyhedron of Conway and Guy in [20]. Recent solutions of two problems on monostable polyhedra [30] by the first-named author, proposed by Conway and Guy [20] in 1969 are also worth mentioning here. As for general convex bodies, a result of Domokos, Papadopoulos and Ruina [12] states that a nondegenerate convex body in the Euclidean plane has at least four equilibria, whereas the construction of a mono-monostatic convex body in [40] by Várkonyi and Domokos, called Gömböc, shows that the same statement does not hold for $3$-dimensional convex bodies. We note that the latter body, the existence of which was conjectured by Arnold [41], is well known outside the mathematics and physics community and has found applications in many other disciplines [^1] For example, it was chosen to symbolize monetary stability by the National Bank of Hungary, and also as the insignia for the Steven Smale Prize by the Society for the Foundations of Computational Mathematics. For more information on the mathematical aspects of centroids and equilibria, the interested reader is referred to the recent survey [27].
@@
In the past century, numerous attempts have been made to generalize the notions of centroids to non-Euclidean geometries, mostly involving spherical and hyperbolic spaces. In particular, centroids of spherical triangles have been investigated by Fog [18] and Fabricius-Bjerre [16] in the 1940s. Their definition uses the model of spherical space as a hypersurface embedded in a Euclidean space (for such an approach, see also [7]). In 1993, Gal’perin [19] gave a thorough investigation of possible definitions of centroids of systems of points in spaces of constant curvature, most importantly in spherical and hyperbolic space. His results were recently generalized for $k$-dimensional manifolds in these spaces in [9]. A recent paper of Besau et al. [5] independently defines the centroid of a spherical set in the same way.
@@
The main goal of this paper is to extend the above investigation to equilibrium properties of convex bodies in non-Euclidean spaces, more specifically in spherical, hyperbolic and normed spaces. In particular, we investigate how the results in [12] and [40] can be generalized for convex bodies in these spaces. During this investigation, we always imagine a $d$-dimensional normed space to be equipped with an underlying Euclidean space.
@@
[…]
In Theorem 1.2, for $d\geq 3$, we call a $d$-dimensional normed space rotationally symmetric if there is a (d2)(d-2)-dimensional linear subspace $L$ in it such that the norm is invariant under any rotation around $L$.
@@ 2.1. Centroids in non-Euclidean spaces
where $d_{E}^{s}(H,x)$ denotes the signed Euclidean distance of $H$ and $x$; the sign is determined by the orientation of $H$. This formula is a consequence of the Euclidean rule of the lever for convex bodies, and the integral in (2.2), denoted by $M_{H}(K)$, is called the first moment of $K$ with respect to $H$.
@@
where $d_{S}^{s}(H,x)$ denotes the signed spherical distance of $H$ and $x$; the sign is determined by the orientation of $H$. Here the integral in (2.4), denoted by $M_{H}(K)$, is called the first moment of $K$ with respect to $H$.
@@
where $d_{H}^{s}(H,x)$ denotes the signed hyperbolic distance of $H$ and $x$; the sign is determined by the orientation of $H$. Here the integral in (2.6), denoted by $M_{H}(K)$, is called the first moment of $K$ with respect to $H$.
@@
    By a classical theorem of Haar, every bounded, translation invariant Borel measure on $\mathbb{R}^{d}$ is a scalar multiple of $d$-dimensional Lebesgue measure. In other words, every ‘meaningful’ definition of volume in $\mathbb{M}^{d}$ is a scalar multiple of $d$-dimensional Euclidean volume (see e.g. [2, 4, 29]).
@@ 2.2. Equilibria in non-Euclidean spaces
Consider a convex body $K$ in the Euclidean space $\mathbb{R}^{d}$. If the point $q\in\operatorname{bd}(K)$ has the property that the hyperplane through $q$ and orthogonal to qc(K)q-c(K) supports $K$, then we say that $q$ is an equilibrium point of $K$. It is well known (see e.g. [10, 11, 30]) that if $K$ is smooth, then the equilibrium points of $K$ coincide with the critical points of the Euclidean distance function $x\mapsto d_{E}(x,c(K))$, measured from $c(K)$ and restricted to $\operatorname{bd}(K)$. We remark that a convex body $K\subset\mathbb{R}^{3}$ is called smooth if for any boundary point $x$ of $K$ there is a unique supporting hyperplane of $K$ at $q$; this property coincides with the property that $\operatorname{bd}(K)$ is a $C^{1}$-class submanifold of $\mathbb{R}^{d}$ (cf. [37]).
@@
Let $K$ be a nondegenerate convex body in $\mathbb{R}^{d}$ with $d=2$ or $d=3$.
@@
By Corollary 2.8, to determine the numbers of the different types of equilibrium points of a nondegenerate convex body in $\mathbb{R}^{2}$, it is sufficient to know the number of its stable points, and in $\mathbb{R}^{3}$ the numbers of its stable and unstable points. Thus, every nondegenerate convex body in $\mathbb{R}^{2}$ belongs to a class $\{S\}_{E}$ consisting of the bodies having $S$ stable and $S$ unstable points, and in $\mathbb{R}^{3}$ to a class $\{S,U\}_{E}$ consisting of the bodies having $S$ stable, $U$ unstable and S+U2S+U-2 saddle-type points. We remark that the global minimizers and maximizers of the Euclidean distance function are stable and unstable points of the body $K$, and thus, the quantities $S,U$ above are positive integers.
@@
For a convex body $K$ in a normed space we follow the same approach. More specifically, if $K$ is a convex body in the normed space $\mathbb{M}^{d}$ and $q\in\operatorname{bd}(K)$, we say that $q$ is an equilibrium point of $K$ if there is a hyperplane $H$ supporting $K$ at $q$ that is Birkhoff orthogonal to the segment $[c(K),q]$. We recall that this is equivalent to the property that a homothetic copy of the unit ball $M$ of $\mathbb{M}^{d}$, centered at $c(K)$, touches $K$. Thus, if both $K$ and $M$ are smooth, then the equilibrium points of $K$ coincide with the critical points of the normed distance function $x\mapsto d_{M}(c(K),x)$ restricted to $\operatorname{bd}(K)$. In the paper we discuss the equilibrium points of convex bodies in planes with strictly convex, smooth norms: in this case for any hyperplane $H$ and point $x\notin H$ there is a unique segment $[x,q]$ with $q\in H$ to which it is Birkoff orthogonal, and for every line $[x,q]$ there is a unique hyperplane through $q$ that is Birkhoff orthogonal to it.
@@
Assume that $x$ is a local minimum point of $\rho_{K}$, and let $D$ be the disk centered at $c(K)$ and satisfying $x\in\operatorname{bd}(D)$. Then, within a neighborhood $V$ of the corresponding point $q_{x}$ of $\operatorname{bd}(K)$, $K$ contains the set $V\cap D$. Since $D$ is smooth and $K$ is convex, the unique supporting line of $D$ at $q_{x}$ supports $K$ as well. Thus, $q_{x}$ is an equilibrium point of $K$. It follows similarly that for every local maximum point $x$ of $\rho_{K}$, the corresponding point $q_{x}$ of $\operatorname{bd}(K)$ is an equilibrium point. Hence, it follows that $\rho_{K}$ is not constant on any nondegenerate arc of $S$, as otherwise $K$ has infinitely many equilibrium points.
@@ 4.1. The case $\mathbb{V}^{3}=\mathbb{S}^{3}$
Consider the unit sphere $\mathbb{S}^{3}$ of $\mathbb{R}^{4}$ centered at the origin $o$. Set the point $v_{N}=(0,0,0,1)$; we call this point the North Pole of $\mathbb{S}^{3}$, and the set of points of $\mathbb{S}^{3}$ with positive $x_{4}$-coordinates the Northern Hemisphere $N^{3}$ of $\mathbb{S}^{3}$. Let $H_{N}$ be the hyperplane of $\mathbb{R}^{4}$ with equation $\{x_{4}=1\}$. The gnomonic projection of $\mathbb{S}^{3}$ onto $H_{N}$ is defined as the map $g_{N}:N^{3}\to H_{N}$, where the image $g_{N}(x)$ of $x\in N^{3}$ is the intersection of $H_{N}$ and the halfline $\{\lambda x:\lambda>0\}$. Note that $g_{N}$ is a bijection between $N^{3}$ and $H_{N}$ with the property that $X\subset N^{3}$ is a greatcircle arc in $N^{3}$ iff its image $g_{N}(X)$ is a Euclidean segment in $H_{N}$. Thus, for any spherically convex body $K\subset N^{3}$, $g_{N}(K)$ is a convex body in $H_{N}$, regarded as a $3$-dimensional Euclidean space, and vice versa: if $L$ is a (Euclidean) convex body in $H_{N}$, then $g^{-1}(L)$ is a (spherical) convex body in $N^{3}\subset\mathbb{S}^{3}$.
@@
$g_{N}^{-1}(x)=\left(\frac{x_{1}}{\sqrt{1+|x|_{N}^{2}}},\frac{x_{2}}{\sqrt{1+|x|_{N}^{2}}},\frac{x_{3}}{\sqrt{1+|x|_{N}^{2}}},\frac{1}{\sqrt{1+|x|_{N}^{2}}}\right).$
@@
For any $0<c\leq 1$, we have $f_{c}\left(\frac{\pi}{2}\right)=g_{c}\left(\frac{\pi}{2}\right)=\frac{\pi}{2}$, $f_{c}\left(-\frac{\pi}{2}\right)=g_{c}\left(-\frac{\pi}{2}\right)=-\frac{\pi}{2}$.
@@
(4.4) $a_{c}(\theta,\varphi)=\frac{\cos^{2}\varphi\cos^{2}f_{c}(\theta)}{\cos^{2}\varphi\cos^{2}f_{c}(\theta)+\sin^{2}\varphi\cos^{2}g_{c}(\theta)},$
as illustrated by Fig. 2. We observe that if $\cos\varphi\neq 0$ (i.e. if $\varphi$ is not of the form $\frac{\pi}{2}+k\pi$ for some integer $k$), then $a_{c}(\theta,\varphi)=\frac{1}{1+\tan^{2}\varphi\frac{\cos^{2}g_{c}(\theta)}{\cos^{2}f_{c}(\theta)}}$.
Figure 2. Diagram of $a_{c}(\theta,\varphi)$ versus $\varphi$ for two different values of $c$. If $c=1$ then $a_{c}=\cos(2\varphi)$ for all values of $\theta$ (left panel), whereas for other values of $c$ and $\theta\neq 0$, it varies. The right panel shows the values of $a_{c}(\theta,\varphi)$ for $c=0.1$, with $\theta=-0.45\pi$ (continuous line), $\theta=-0.225\theta$ (long dashed line), $\theta=0$ (dashed line), $\theta=0.225\pi$ (dashed-dotted line), and $\theta=0.45\pi$ (dotted line).
@@
and remark that by definition, for any value of the parameters we have $0\leq a_{c}(\theta,\varphi)\leq 1$. Thus, $\rho_{c}(u)$ is a convex combination of $f_{c}(\theta)$ and $g_{c}(\theta)$. This, by Lemma 4.3, implies that $1\leq\rho_{c}(u)\leq 1$ and $R(1-d)\leq R_{c,d}(u)\leq R(1+d)$, implying that $K(c,d)$ exists for all $0\leq d<1$.
@@
Next, we prove that the properties described in (A)-(B) are satisfied for $K(c,d)$ for all $0<d<1$ and $0<c\leq 1$. To show them, we recall [11, Lemmas 3, 4]; we state them using the notation introduced in this subsection.
@@
Now we prove that suitable elements of $\mathcal{F}$ satisfy the properties (C)-(E). For this purpose, we set some arbitrary small positive constant $\varepsilon>0$, and observe that the definition of $\rho_{c}$ implies that $\max\{\rho_{c}(u):u\in S\}=1$, and hence, if $0<d<\varepsilon$, then $d_{H}(K_{n}(c,d),RB)<R\varepsilon$. Thus, the assertion of Theorem 1.2 for $\mathbb{S}^{3}$ readily follows from Lemmas 4.7 and 4.8.
@@
There exist constants $0<c_{1}<c_{2}<1$, $0<d_{0}<1$ and a function $F:(0,d_{0})\to[c_{1},c_{2}]$ such that for any $d\in(0,d_{0})$ the (spherical) centroid of $K(F(d),d)$ is $v_{N}$.
@@
We note that the integral in (4.9) is also computed in [11], where the authors incorrectly stated that its value is $2.3168\ldots$; nevertheless, this typo in [11] does not influence the correctness of their proof.
Next, we deal with Property (C). To do it, following [37] we denote by $\mathcal{C}^{2}_{+}$ the family of convex bodies whose boundary is $C^{2}$-class differentiable at every point, and has strictly positive Gaussian curvature everywhere.